# Dynamics of Meta-learning Representation in the Teacher-student Scenario

### v1

@@ III.1 The Meta-Teacher and Task-specific Teachers
To address meta-learning across various tasks, we present a meta-teacher model designed to generate multiple specific teachers tailored to different tasks. These specific teachers share similar representations with the meta-teacher but differ from one another based on the supervised learning tasks built on top of the representation. Specifically, both the meta-teacher and the specific teachers are two-layer neural networks, having $N$ input units, $M$ hidden units and one output unit. The meta-teacher has input-to-hidden weights $\mathbf{B}\in\mathbb{R}^{M\times N}$ and hidden-to-output weights $\boldsymbol{u}^{0}\in\mathbb{R}^{M}$. For the task-specific teachers, we will firstly assume that their input-to-hidden weights are the same and copied from the meta-teacher, having values $\mathbf{B}$. We will relax this constraint in a later section. Conversely, the hidden-to-output weights $\boldsymbol{u}^{t}\in\mathbb{R}^{M}$ are specific to the $t$-th task, assumed to be drawn from a multivariate Gaussian distribution $\boldsymbol{u}^{t}\sim\mathcal{N}(0,I_{M})$. In this way, the weights $\{\boldsymbol{u}^{t}\}$ are completely independent among different tasks. A less challenging scenario for the meta-learner is to create some level of similarity among different $\{\boldsymbol{u}^{t}\}$, but we will primarily focus on the regime of uncorrelated $\{\boldsymbol{u}^{t}\}$ to simplify the picture. The teacher-task generation process is illustrated in Fig 1(a). In the notation of Sec. II, task $\mathcal{T}_{t}$ is fixed by the parameters $\{\mathbf{B},\boldsymbol{u}^{t}\}$ of the $t$-th teacher network.
Figure 1: The framework of the teacher-student analysis for meta-learning under consideration. All models are two-layer neural networks. (a) The meta-teacher generates different task-specific teachers, which share the same input-to-hidden weights with the meta-teacher (denoted as $\mathbf{B}$) but have different hidden-to-output weights (denoted as $\boldsymbol{u}^{t}$). (b) The meta-training and testing processes. The meta-learner is meta-trained on different tasks by using the FO-ANIL algorithm. Each specific learner inherits the input-to-hidden weights $\mathbf{J}$ from the meta-learner, but maintains its own hidden-to-output weights (denoted as $\boldsymbol{w}^{t}$) which are task specific. The meta-testing is evaluated by its ability to guide a new learner to solve an unseen task $\mathcal{T}_{\text{new}}$ (through an inner-loop optimization).
The training dataset for the $t$-th task can be prepared by separately feeding $P$ inputs $\{\boldsymbol{\xi}^{t,\mu}\}_{\mu=1}^{P}$ to the $t$-th specific teacher network and collecting the outputs
@@
where $\boldsymbol{B}_{m}\in\mathbb{R}^{1\times N}$ is the $m$-th row of the matrix $\mathbf{B}$ (or the weight vector of the $m$-th hidden unit), and $g(\cdot)$ is the activation function. The factor $1/\sqrt{M}$ in Eq. (7) was introduced to ensure that the outputs remain of the same order of magnitude across networks with different hidden layer widths. The validation or test dataset can be obtained similarly. We primarily consider the error activation function $g(x)=\operatorname{erf}(x/\sqrt{2})$ for analytical tractability, and take the high dimension limit of the input $N\to\infty$. Moreover, we concern about only a finite number of hidden units in all the teacher networks (i.e., $M\sim\mathcal{O}(1)$). Another interesting case involves networks of infinite width with $M\to\infty$ (as in [32, 33]), but analyzing the learning dynamics in such cases requires different assumptions and theoretical tools [34, 35, 36], which are beyond the scope of this study.
@@ III.2 The Meta-learner and Task-specific Students
The meta-learner and the student networks for specific tasks are also assumed to be two-layer neural networks, having $N$ input units, $K$ hidden units and one output unit. We also consider that $K$ is finite (i.e., $K\sim\mathcal{O}(1)$), which is not necessary the same as $M$. The meta-learner has input-to-hidden weights $\mathbf{J}\in\mathbb{R}^{K\times N}$ and hidden-to-output weights $\boldsymbol{w}^{0}\in\mathbb{R}^{1\times K}$. For the $t$-th task, the student network copies the input-to-hidden weights $\mathbf{J}$ from the meta-learner, saving the effort to learn the representation mapping; but it maintains its own hidden-to-output weights $\boldsymbol{w}^{t}\in\mathbb{R}^{1\times K}$. Such a set-up is in alignment with the ANIL algorithm, where the representation mapping is not fine-tuned when the $t$-student is adapted to solve the $t$-task.
Feeding the inputs $\{\boldsymbol{\xi}^{t,\mu}\}_{\mu=1}^{P}$ to the $t$-th student network gives rise to its output predictions
@@
The objective of the $t$-th student is to adapt the weights $\boldsymbol{w}^{t}$ such that the above outputs $\{\tilde{\sigma}^{t,\mu}\}_{\mu=1}^{P}$ get close to the labels $\{\sigma^{t,\mu}\}_{\mu=1}^{P}$ given in Eq. (7) in order to solve the task $\mathcal{T}_{t}$.
@@ III.3 Meta-training and Meta-testing
Referring to the notations in Sec. II, we identify that the model parameters of the $t$-th student are $\boldsymbol{\theta}_{t}=\{\mathbf{J},\boldsymbol{w}^{t}\}$. In the MAML algorithm, both $\mathbf{J}$ and $\boldsymbol{w}^{0}$ are adapted in the inner-loop optimization when solving the specific task $\mathcal{T}_{t}$. On the other hand, in the ANIL algorithm, the input-to-hidden weights $\mathbf{J}$ which maps the inputs to the latent representations are fixed in the inner-loop optimization, as we have exploited in Sec. III.2 for the set-up of teacher-student analysis. Furthermore, the FO-ANIL algorithm neglects any possible second derivatives in the vanilla ANIL algorithm to simplify the computations.
We consider the case that the tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{t},\dots\}$ drawn from the distribution $\mathcal{P}(\mathcal{T})$ arrive in a streaming, online fashion where $t$ can be viewed as a time index. Each task comes with its own training and validating data sets, generated by the meta-teacher framework described above using independent inputs $\boldsymbol{\xi}^{t,\mu}$. For the task $\mathcal{T}_{t}$, the $t$-th student network receives a small training set $D_{\text{train}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\mu},\sigma^{t,\mu})\}_{\mu=1}^{P}$. The training loss for the inner-loop optimization is defined as
@@
where the pre-factor $\frac{1}{2}$ is introduced for notational convenience. The FO-ANIL algorithm dictates that the $t$-th student network slightly adapts the hidden-to-output weights by lowering this training loss starting from the parameters of the meta-learner
$\boldsymbol{w}^{t}=\boldsymbol{w}^{0}-\eta_{w}\nabla_{\boldsymbol{w}}\big{[}\ell(\boldsymbol{w}\mid\mathbf{J},D_{\text{train}}^{\mathcal{T}_{t}})\big{]}\big{|}_{\boldsymbol{w}=\boldsymbol{w}^{0}},$ (10)
@@
In this special case, the meta-learner has no need to maintain the parameters $\boldsymbol{w}^{0}$, and the meta-training resorts to learning a better representation mapping $\mathbf{J}$. To achieve this, we firstly define the validation loss for the outer optimization using the validation dataset $\mathcal{D}_{\text{val}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\nu},\sigma^{t,\nu})\}_{\nu=1}^{V}$
@@
$\boldsymbol{J}_{k}^{t}=\boldsymbol{J}_{k}^{t-1}-\frac{\eta_{J}}{N}\big{[}\nabla_{\boldsymbol{J}_{k}}\ell(\mathbf{J}\mid D_{\text{val}}^{\mathcal{T}_{t}})\big{]}\big{|}_{\boldsymbol{J}_{k}=\boldsymbol{J}_{k}^{t-1}},$ (13)
@@
Suppose that at a certain time, the meta-learner obtains a meta-representation mapping $\mathbf{J}$, where we would like to evaluate its performance. This can be done by firstly considering a new task $\mathcal{T}_{\text{new}}$ drawn from $\mathcal{P}(\mathcal{T})$, which comes with a small training set $D_{\text{train}}^{\mathcal{T}_{\text{new}}}=\{(\boldsymbol{\xi}^{{\text{new},\mu}},\sigma^{{\text{new},\mu}})\}_{\mu=1}^{P}$ generated by the new task’s teacher network. A new student network $\mathcal{L}_{\text{new}}$ will be assigned for this task; it will adapt its hidden-to-output weights according to Eq. (11), yielding $\boldsymbol{w}^{\text{new}}$. The performance of this new student is measured by the generalization error when predicting on an unseen test data point $(\boldsymbol{\xi}^{\text{new},0},\sigma^{\text{new},0})$, average over the input distribution of $\boldsymbol{\xi}^{\text{new},0}$
@@ III.4 Macroscopic Order Parameters
We denote the pre-activation of $k$-th hidden unit in the meta-learner’s network as $x_{k}^{t,\mu}=\boldsymbol{J}_{k}^{\top}\cdot\boldsymbol{\xi}^{t,\mu}$, and the pre-activation of the $m$-th hidden unit in the meta-teacher’s network as $y_{k}^{t,\mu}=\boldsymbol{B}_{k}^{\top}\cdot\boldsymbol{\xi}^{t,\mu}$. Recall that the elements of $\boldsymbol{\xi}^{t,\mu}$ are independent Gaussian random variables with zero mean and unit variance. In the limit where the input dimension is large, i.e., $N\to\infty$, the central limit theorem dictates that the meta-learner’s and the meta-teacher’s pre-activations follow a joint Gaussian distribution, where the means are $\langle x_{k}^{t,\mu}\rangle=\langle y_{m}^{t,\mu}\rangle=0$ and the covariance matrix is specified by
@@
More explicitly, the local fields $(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})\in\mathbb{R}^{K+M}$ obeys the following Gaussian distribution
$\!\!\mathcal{P}(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})$ $=\frac{\exp\{-\frac{1}{2}(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})^{\top}C^{-1}(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})\}}{\sqrt{(2\pi)^{K+M}|C|}},$ (19)
@@
We recall that the label $\sigma^{t,\mu}$ is generated by the $t$-th teacher network in Eq. (7), so the updated hidden-to-output weights $w^{t}_{k}$ of the $t$-th student network admit the following expression
@@
where the randomness comes from the training inputs $\{\boldsymbol{\xi}^{t,\mu}\}_{\mu=1}^{P}$ and the task vector $\boldsymbol{u}^{t}$ for task $\mathcal{T}_{t}$.
@@
×$\quad\;\;\times w_{k}^{t}g^{\prime}(x_{k}^{t,\nu}).$ (25)
where the randomness of the pre-activation $\{x_{k}^{t,\nu},y_{m}^{t,\nu}\}$ comes from the validation inputs $\{\boldsymbol{\xi}^{t,\nu}\}_{\nu=1}^{V}$.
@@
The evolution of the representation mapping in Eq. (24) involves high-dimensional microscopic dynamical variables $\boldsymbol{J}_{k}^{t}$. It can be projected onto the low dimensional space of order parameters by multiplying both sides of Eq. (24) with $\boldsymbol{B}_{n}^{t}$ and $\boldsymbol{J}_{l}^{t}$, leading to the evolution of two set of order parameters
@@
Traditional statistical mechanics analysis of online learning arrives at a similar set of dynamical equations of order parameters, and proceeds to examine their deterministic evolution assuming the self-averaging property of the order parameters. For our meta-learning problem, it is a bit more tricky to do that, as the right hand sides of the Eqs. (26) and ( 27) have three sources of randomness, (i) from the local fields $(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})$ through the stochastic inputs $\{\boldsymbol{\xi}^{t,\mu}\}$, (ii) from the local fields $(\boldsymbol{x}^{t,\nu},\boldsymbol{y}^{t,\nu})$ through the stochastic inputs $\{\boldsymbol{\xi}^{t,\nu}\}$, (iii) from the choice of the $t$-th teacher’s task vector $\boldsymbol{u}^{t}$. In Appendix B, we provide a heuristic argument to show that when $P$ and $V$ are large, the dynamics of the order parameters $\{R_{kn},Q_{kl}\}$ admit the self-averaging property, i.e., a realization of their random trajectory stays close to their average values. Nevertheless, the values of $P$ and $V$ can still be significantly smaller than those required for successful learning in single-task settings.
@@
$\frac{\mathrm{d}Q_{kl}}{\mathrm{d}\alpha}$ $=\frac{\eta_{J}}{V\sqrt{K}}\sum_{\nu=1}^{V}\big{[}\big{\langle}h_{k}^{\nu}x_{l}^{\nu}\big{\rangle}+\big{\langle}h_{l}^{\nu}x_{k}^{\nu}\big{\rangle}\big{]}+\frac{\eta_{J}^{2}}{V^{2}K}\big{\langle}h_{k}^{\nu}h_{l}^{\nu}\big{\rangle}.$ (29)

### v2

@@ III.1 The meta-teacher and task-specific teachers
To address meta-learning across various tasks, we present a meta-teacher model designed to generate multiple specific teachers tailored to different tasks. These specific teachers share similar representations with the meta-teacher but differ from one another based on the supervised learning tasks built on top of the representation. Specifically, both the meta-teacher and the specific teachers are two-layer neural networks, having $N$ input units, $M$ hidden units and one output unit. The meta-teacher has input-to-hidden weights $\mathbf{B}\in\mathbb{R}^{M\times N}$ and hidden-to-output weights $\boldsymbol{u}^{0}\in\mathbb{R}^{M}$. In this study, we set $\boldsymbol{u}^{0}$ to zero, indicating that the meta-teacher does not impose a preferred function but instead provides a shared representation mapping.
For the task-specific teachers, we will first assume that their input-to-hidden weights are the same and copied from the meta-teacher, having values $\mathbf{B}$. We will relax this constraint in a later section. Conversely, the hidden-to-output weights $\boldsymbol{u}^{t}\in\mathbb{R}^{M}$ are specific to the $t$th task, assumed to be drawn from a multivariate Gaussian distribution $\boldsymbol{u}^{t}\sim\mathcal{N}(0,I_{M})$. In this way, the weights $\{\boldsymbol{u}^{t}\}$ are completely independent among different tasks, which in fact creates a challenging scenario for the meta-learner. A less challenging scenario for the meta-learner is to create some level of similarity among different $\{\boldsymbol{u}^{t}\}$, but we will primarily focus on the regime of uncorrelated $\{\boldsymbol{u}^{t}\}$ to simplify the picture. The teacher-task generation process is illustrated in Fig 1(a). In the notation of Sec. II, task $\mathcal{T}_{t}$ is fixed by the parameters $\{\mathbf{B},\boldsymbol{u}^{t}\}$ of the $t$th teacher network.
Figure 1: The framework of the teacher-student analysis for meta-learning under consideration. All models are two-layer neural networks. (a) The meta-teacher generates different task-specific teachers, which share the same input-to-hidden weights with the meta-teacher (denoted as $\mathbf{B}$) but have different hidden-to-output weights (denoted as $\boldsymbol{u}^{t}$). (b) The meta-training and testing processes. The meta-learner is meta-trained on different tasks using the FO-ANIL algorithm. Each specific learner inherits the input-to-hidden weights $\mathbf{J}$ from the meta-learner, but maintains its own hidden-to-output weights (denoted as $\boldsymbol{w}^{t}$) which are task specific. The meta-testing is evaluated by its ability to guide a new learner to solve an unseen task $\mathcal{T}_{\text{new}}$ (through an inner-loop optimization).
The training dataset for the $t$th task can be prepared by separately feeding $P$ inputs $\{\boldsymbol{\xi}^{t,\mu}\}_{\mu=1}^{P}$ to the $t$th specific teacher network and collecting the outputs
@@
where $\boldsymbol{B}_{m}\in\mathbb{R}^{1\times N}$ is the $m$th row of the matrix $\mathbf{B}$ (or the weight vector of the $m$th hidden unit), and $g(\cdot)$ is the activation function. The factor $1/\sqrt{M}$ in Eq. (7) was introduced to ensure that the outputs remain of the same order of magnitude across networks with different hidden layer widths. The validation or test dataset can be obtained similarly. We primarily consider the error activation function $g(x)=\operatorname{erf}(x/\sqrt{2})$ for analytical tractability, and take the high-dimension limit of the input $N\to\infty$. Moreover, we concern about only a finite number of hidden units in all the teacher networks [i.e., $M\sim\mathit{O}(1)$]. Another interesting case involves networks of infinite width with $M\to\infty$ (as in [32, 33]), but analyzing the learning dynamics in such cases requires different assumptions and theoretical tools [34, 35, 36], which are beyond the scope of this study.
@@ III.2 The meta-learner and task-specific students
The meta-learner and the student networks for specific tasks are also assumed to be two-layer neural networks, having $N$ input units, $K$ hidden units and one output unit. We also consider that $K$ is finite [i.e., $K\sim\mathit{O}(1)$], which is not necessarily the same as $M$. The meta-learner has input-to-hidden weights $\mathbf{J}\in\mathbb{R}^{K\times N}$ and hidden-to-output weights $\boldsymbol{w}^{0}\in\mathbb{R}^{1\times K}$. For the $t$th task, the student network copies the input-to-hidden weights $\mathbf{J}$ from the meta-learner, saving the effort to learn the representation mapping, but it maintains its own hidden-to-output weights $\boldsymbol{w}^{t}\in\mathbb{R}^{1\times K}$. Such a setup is in alignment with the ANIL algorithm, where the representation mapping is not fine-tuned when the $t$-student is adapted to solve the $t$-task.
Feeding the inputs $\{\boldsymbol{\xi}^{t,\mu}\}_{\mu=1}^{P}$ to the $t$th student network gives rise to its output predictions
@@
The objective of the $t$th student is to adapt the weights $\boldsymbol{w}^{t}$ such that the above outputs $\{\tilde{\sigma}^{t,\mu}\}_{\mu=1}^{P}$ get close to the labels $\{\sigma^{t,\mu}\}_{\mu=1}^{P}$ given in Eq. (7) in order to solve the task $\mathcal{T}_{t}$.
@@ III.3 Meta-training and meta-testing
Referring to the notations in Sec. II, we identify that the model parameters of the $t$th student are $\boldsymbol{\theta}_{t}=\{\mathbf{J},\boldsymbol{w}^{t}\}$. In the MAML algorithm, both $\mathbf{J}$ and $\boldsymbol{w}^{0}$ are adapted in the inner-loop optimization when solving the specific task $\mathcal{T}_{t}$. On the other hand, in the ANIL algorithm, the input-to-hidden weights $\mathbf{J}$ which map the inputs to the latent representations are fixed in the inner-loop optimization, as we have exploited in Sec. III.2 for the setup of teacher-student analysis. Furthermore, the FO-ANIL algorithm neglects any possible second derivatives in the vanilla ANIL algorithm to simplify the computations.
We consider the case that the tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{t},\dots\}$ drawn from the distribution $\mathcal{P}(\mathcal{T})$ arrive in a streaming, online fashion where $t$ can be viewed as a time index. Each task comes with its own training and validating data sets, generated by the meta-teacher framework described above using independent inputs $\boldsymbol{\xi}^{t,\mu}$. For the task $\mathcal{T}_{t}$, the $t$th student network receives a small training set $D_{\text{train}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\mu},\sigma^{t,\mu})\}_{\mu=1}^{P}$. The training loss for the inner-loop optimization is defined as
@@
where the prefactor $\frac{1}{2}$ is introduced for notational convenience. The FO-ANIL algorithm dictates that the $t$th student network slightly adapts the hidden-to-output weights by lowering this training loss starting from the parameters of the meta-learner
$\boldsymbol{w}^{t}=\boldsymbol{w}^{0}-\eta_{w}\mathit{\nabla}_{\boldsymbol{w}}\big{[}\ell(\boldsymbol{w}\mid\mathbf{J},D_{\text{train}}^{\mathcal{T}_{t}})\big{]}\big{|}_{\boldsymbol{w}=\boldsymbol{w}^{0}},$ (10)
@@
In this special case, the meta-learner has no need to maintain the parameters $\boldsymbol{w}^{0}$, and the meta-training resorts to learning a better representation mapping $\mathbf{J}$. To achieve this, we first define the validation loss for the outer optimization using the validation dataset $D_{\text{val}}^{\mathcal{T}_{t}}=\{(\boldsymbol{\xi}^{t,\nu},\sigma^{t,\nu})\}_{\nu=1}^{V}$
@@
$\boldsymbol{J}_{k}^{t}=\boldsymbol{J}_{k}^{t-1}-\frac{\eta_{J}}{N}\big{[}\mathit{\nabla}_{\boldsymbol{J}_{k}}\ell(\mathbf{J}\mid D_{\text{val}}^{\mathcal{T}_{t}})\big{]}\big{|}_{\boldsymbol{J}_{k}=\boldsymbol{J}_{k}^{t-1}},$ (13)
@@
Suppose that at a certain time, the meta-learner obtains a meta-representation mapping $\mathbf{J}$, where we would like to evaluate its performance. This can be done by first considering a new task $\mathcal{T}_{\text{new}}$ drawn from $\mathcal{P}(\mathcal{T})$, which comes with a small training set $D_{\text{train}}^{\mathcal{T}_{\text{new}}}=\{(\boldsymbol{\xi}^{{\text{new},\mu}},\sigma^{{\text{new},\mu}})\}_{\mu=1}^{P}$ generated by the new task’s teacher network. A new student network $\mathcal{L}_{\text{new}}$ will be assigned for this task; it will adapt its hidden-to-output weights according to Eq. (11), yielding $\boldsymbol{w}^{\text{new}}$. The performance of this new student is measured by the generalization error when predicting an unseen test data point $(\boldsymbol{\xi}^{\text{new},0},\sigma^{\text{new},0})$, average over the input distribution of $\boldsymbol{\xi}^{\text{new},0}$
@@ III.4 Macroscopic order parameters
We denote the preactivation of $k$th hidden unit in the meta-learner’s network as $x_{k}^{t,\mu}=\boldsymbol{J}_{k}^{\top}\cdot\boldsymbol{\xi}^{t,\mu}$, and the preactivation of the $m$th hidden unit in the meta-teacher’s network as $y_{k}^{t,\mu}=\boldsymbol{B}_{k}^{\top}\cdot\boldsymbol{\xi}^{t,\mu}$. Recall that the elements of $\boldsymbol{\xi}^{t,\mu}$ are independent Gaussian random variables with zero mean and unit variance. In the limit where the input dimension is large, i.e., $N\to\infty$, the central limit theorem dictates that the meta-learner’s and the meta-teacher’s preactivations follow a joint Gaussian distribution, where the means are $\langle x_{k}^{t,\mu}\rangle=\langle y_{m}^{t,\mu}\rangle=0$ and the covariance matrix is specified by
@@
More explicitly, the local fields $(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})\in\mathbb{R}^{K+M}$ obey the following Gaussian distribution
$\!\!\!\mathcal{P}(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})$ $=\frac{\exp\left[-\frac{1}{2}(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})^{\top}C^{-1}(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})\right]}{\sqrt{(2\pi)^{K+M}|C|}},$ (19)
@@
We recall that the label $\sigma^{t,\mu}$ is generated by the $t$th teacher network in Eq. (7), so the updated hidden-to-output weights $w^{t}_{k}$ of the $t$th student network admit the following expression:
@@
where the randomness comes from the training inputs $\{\boldsymbol{\xi}^{t,\mu}\}_{\mu=1}^{P}$ and the task vector $\boldsymbol{u}^{t}$ for task $\mathcal{T}_{t}$. We note that the dynamics of $w^{t}_{k}$ are not maintained, as $w^{t}_{k}$ is not derived from $w^{t-1}_{k}$. This is in stark contrast to the behavior of hidden-to-output weights in the single-task learning scenario analyzed in Ref. [40].
@@
×$\quad\;\;\times w_{k}^{t}g^{\prime}(x_{k}^{t,\nu}),$ (25)
where the randomness of the preactivation $\{x_{k}^{t,\nu},y_{m}^{t,\nu}\}$ comes from the validation inputs $\{\boldsymbol{\xi}^{t,\nu}\}_{\nu=1}^{V}$.
@@
The evolution of the representation mapping in Eq. (24) involves high-dimensional microscopic dynamical variables $\boldsymbol{J}_{k}^{t}$. It can be projected onto the low-dimensional space of order parameters by multiplying both sides of Eq. (24) with $\boldsymbol{B}_{n}^{t}$ and $\boldsymbol{J}_{l}^{t}$, leading to the evolution of two sets of order parameters,
@@
Traditional statistical mechanics analysis of online learning arrives at a similar set of dynamical equations of order parameters, and proceeds to examine their deterministic evolution assuming the self-averaging property of the order parameters. For our meta-learning problem, it is a bit more tricky to do that, as the right-hand sides of the Eqs. (26) and ( 27) have three sources of randomness, (i) from the local fields $(\boldsymbol{x}^{t,\mu},\boldsymbol{y}^{t,\mu})$ through the stochastic inputs $\{\boldsymbol{\xi}^{t,\mu}\}$, (ii) from the local fields $(\boldsymbol{x}^{t,\nu},\boldsymbol{y}^{t,\nu})$ through the stochastic inputs $\{\boldsymbol{\xi}^{t,\nu}\}$, and (iii) from the choice of the $t$th teacher’s task vector $\boldsymbol{u}^{t}$. In Appendix C, we provide a heuristic argument to show that when $P$ and $V$ are large, the dynamics of the order parameters $\{R_{kn},Q_{kl}\}$ admit the self-averaging property, i.e., a realization of their random trajectory stays close to their average values. Nevertheless, the values of $P$ and $V$ can still be significantly smaller than those required for successful learning in single-task settings.
@@
$\frac{\mathrm{d}Q_{kl}}{\mathrm{d}\alpha}$ $=\frac{\eta_{J}}{V\sqrt{K}}\sum_{\nu=1}^{V}\big{[}\big{\langle}h_{k}^{\nu}x_{l}^{\nu}\big{\rangle}+\big{\langle}h_{l}^{\nu}x_{k}^{\nu}\big{\rangle}\big{]}+\frac{\eta_{J}^{2}}{V^{2}K}\big{\langle}h_{k}^{\nu}h_{l}^{\nu}\big{\rangle},$ (29)